A bank pays 10% nominal annual interest on special three-year certificates. What is the effective annual interest rate if interest is compounded
a. Every three months?
b. Daily?
c. Continuously?
1
Expert's answer
2021-09-15T02:15:14-0400

a. Every three months

"r=(1+\frac{0.10}{4})^4-1\\
r=(1.025)^4-1\\
r=0.1038\\
r=10.38\%"


b. Daily

"r=(1+\frac{0.10}{365})^{365}-1\\
r=(1.0757)^{365}-1\\
r=0.0757\\
r=7.57\%"


c. Continuously

"r=e^{0.10}-1\\
r=1.1051-1\\
r=0.1501\\
r=1.501\%"